Money, Banking, and the Federal Reserve System — End of Chapter Problem

What will happen to the money supply under the following circumstances in a checkable–deposits–only system? Calculate the change in the money supply in each circumstance.

a. The required reserve ratio is 25%, and a depositor withdraws $700 from his checkable bank deposit.

Change in money supply: $



b. The required reserve ratio is 5%, and a depositor withdraws $700 from his checkable bank deposit.

Change in money supply: $



c. The required reserve ratio is 20%, and a customer deposits $750 into her checkable bank deposit.

Change in money supply: $



d. The required reserve ratio is 10%, and a customer deposits $600 into her checkable bank deposit.

Change in money supply: $



1
Expert's answer
2021-04-02T07:29:21-0400

Change in money supply = (1/ RR) * monetary base


If there is a deposit than there is an positive increase in reserves


If there is a withdrawal there is a negative or decline in reserves


A.RR = 25% = 0.25


M = 700


Change in money supply ="(\n\n\n\n 1\/0.25)*700 = 2800\n\nB.RR = 5% = 0.05" ( 1/0.25)*700 = 2800


M = 700


Change in money supply = "(1\/0.05)*700 = 14000\n\n\n\nC. RR = 20% = 0.2" (1/0.05)*700 = 14000


C. RR = 20% = 0.2


M = 750


Change in money supply ="(1\/0.2)*750 = 3750\n\n\n\nD. RR = 10% = 0.1" (1/0.2)*750 = 3750


D. RR = 10% = 0.1


M = 600


Change in money supply ="600*(1\/0.1) = 6000"


2) When there is a recessionary gap the central bank will adopt an expansionary monetary policy where it increases the money supply in the market and lower the interest rates. So that there is an increase in aggregate demand.


When there is an expansionary monetary policy there will be an increase in consumer spending , lending and investments. Aggregate demand also increases as aggregate price level increases.